# ADVANCING MATHEMATICS RESEARCH WITH GENERATIVE AI

### v2

@@ Abstract.
The main drawback of using generative AI models for advanced mathematics is that these models are not logical reasoning engines. However, Large Language Models, and their refinements, can pick up on patterns in higher mathematics that are difficult for humans to see. By putting the design of generative AI models to their advantage, mathematicians may use them as powerful interactive assistants that can carry out laborious tasks, generate and debug code, check examples, formulate conjectures and more. We discuss how generative AI models can be used to advance mathematics research. We also discuss their integration with Computer Algebra Systems and formal proof assistants such as Lean.
@@ 1. Introduction
This shortfall is built into the design of LLMs. They are fundamentally statistical, not logical, engines. As next-word predictors, they lack a built-in engine for formal symbolic deduction, meaning their emergent ‘reasoning’, called Natural Language Reasoning, is derived from statistical patterns in their training data.
@@
The AI landscape is rapidly evolving beyond standard LLMs, with Large Reasoning Models (LRMs) and Large Context Models (LCMs) emerging as the next generation of this technology. While built on the foundation of LLMs, these models have distinct architectures and methods of training (see Section 4).
In Section 2, we discuss the mechanics of probabilistic language generation. In Section 3, we discuss how to manage the context window. In Sections  59, we describe the capabilities and limitations of LLMs, the nature of their training data, and their ability to detect geometric patterns in high-dimensional representations of mathematical language. In Sections 1016, we discuss applications and user techniques. This includes examples from combinatorial group theory, a comparison of current models, a guide to prompt engineering, and fact-checking. In Section 17 we discuss the integration of LLMs with Computer Algebra Systems (CAS) and formal proof assistants like Lean.
2. The mechanics of probabilistic language generation
2.1. The LLM as a probabilistic knowledge graph
@@ 2.2. Tokens: the building blocks of language
2.3. How the prompt shapes an LLM’s output
3. A single chat is a ‘context window’
@@
A generative AI model’s short-term memory is its context window, which consists of a single chat session. This window contains the entire conversation history, including uploaded documents and any corrections made by the user. Its capacity is measured in tokens: while a 120K window can hold a novel, models like Gemini 2.5 Pro offer an unprecedented 1 million tokens.
@@
4. Large Reasoning Models and Large Context Models
Large Reasoning Models (LRMs), also known as a Reasoning Language Model (RLMs), are a class of LLMs trained to carry out tasks that require multi-step deduction, critical thinking, and structured problem-solving. LRMs ‘think’ before they respond.
They deliberate, explore various solution paths, evaluate intermediate steps and revise their reasoning.
LRMs integrate frameworks that support reasoning structures, such as ‘chain-of-thought’ and ‘tree-of-thought’ structures into their systems. We give some examples in Section 14.
@@ 4. Large Reasoning Models and Large Context Models
Google’s Gemini models, particularly the 2.5 series, demonstrate both LRM and LCM characteristics. The models also use internal ‘thinking processes’ for improved reasoning.
ChatGPT has a ‘reasoning mode’ in its model OpenAI o3. Reinforcement learning was used to teach OpenAI o3 to ‘think’ before generating answers, using what OpenAI refers to as ‘private chain-of-thought’ processes built into their system prompts. This allows the model to plan ahead and reason through tasks, performing a series of intermediate reasoning steps to assist in solving the problem.
Claude models use ‘extended thinking’ or the ability to ‘think out loud’. As LCMs, they offer substantial context windows
5. Drawbacks of using generative AI in mathematics
The primary drawback of using current AI models for mathematics is that they are probabilistic pattern-matchers, not logical reasoning engines. Their core function is to predict the next most likely word or symbol in a sequence, not to apply deterministic mathematical rules. Generative AI models are experts at reconstructing what a proof should ‘look like’. This limitation leads to several key problems:
Consequently, mathematical output from an LLM has to be carefully scrutinized for logical and symbolic correctness.
6. Mathematical training data for generative AI models
7. Generative AI models simulate computer algebra systems
@@
In addition, the queries may give incorrect answers if the problems are inaccurately posed. The LLM could misunderstand the prompt, have bugs in its code or fail to handle boundary cases.
@@
8. How much do generative AI models ‘know’?
In contrast, Symbolic Deduction is manipulation based on formal rules applied to abstract symbols, as in the operation of a Computer Algebra System. Current generative AI models are built to imitate reasoning based on patterns, but are incapable of performing true symbolic deduction.
The significant advancement in reasoning capability has been achieved by an advanced form of chain-of-thought and tree-of-thought reasoning. These are important tools when using generative AI models for mathematics. We give some examples in Section 14.
9. generative AI models and patterns in higher mathematics
A generative AI model can identify mathematical patterns that are imperceptible to humans. It performs massive scale statistical analysis on the symbolic structure of mathematics itself. This allows it to simulate a mathematician’s intuition. It can find cross-disciplinary correlations and potential analogies that humans may not think of. It is unconstrained by human cognitive biases and the boundaries of human knowledge.
9.1. Advantages of this capability:
A generative AI model sees mathematics as a language with a complex grammar and vocabulary composed of symbols and diagrams. It learns the statistical relationships between the tokens in this language.
Generative AI models don’t ‘see’ symbols and equations. There is a map $\varphi:V\to\mathbb{R}^{d}$ (for $d\leq 13,000$) from the vocabulary $V$ of tokens to a high-dimensional vector space (the embedding space). The dimension $d\sim|V|^{1/4}$ is a model parameter. Relationships in the vocabulary are translated into geometric relationships in $\mathbb{R}^{d}$ [mikolov2013].
The model can detect geometric relationships in $\mathbb{R}^{d}$ that are impossible for humans to visualize and that are not obvious from the symbolic definitions.
The image $\varphi(w)$ of a word $w$ doesn’t have a single location in the embedding space. Its ‘meaning’ is its relationship to all other images of tokens in the space and their locations. Furthermore, this meaning changes with context. The model applies a linear transformation $T$ to $\varphi(w)\in\mathbb{R}^{d}$ to project it onto a subspace relevant to its contextual role. This allows it to isolate a specific feature to look at. It then reviews all isolated facts and uses piecewise linear maps to put these facts together and compile an output.
This ability to perceive and analyze non-obvious geometric patterns allows the AI model to generate novel conjectures and insights when suitably prompted.
9.2. Human-LLM collaborative research
10. Why AI is a useful tool in combinatorial group theory
As an example, we solve the word problem for the word $w=yxy^{-2}x^{-1}$:
We first rewrite $x^{-1}yx=y^{2}$ as $yx=xy^{2}$. Then
${yx}y^{-2}x^{-1}={xy^{2}}y^{-2}x^{-1}=xx^{-1}=1.$
Thus the word $w=yxy^{-2}x^{-1}$ represents the identity element in $G$.
10.1. Solving algorithmic research questions
Algorithmic search problems in group theory, such as the word problem, conjugacy problem (deciding if two words are conjugate in a given group), and triviality problem (deciding if a group presentation presents the trivial group), are undecidable in general. Even in decidable cases, the search space for groups with many generators and intertwined relations is often astronomically large, making it intractable for hand calculation.
However, this is not true symbolic manipulation: generative AI models mimic the process through pattern recognition, so their output must always be verified and corrected.

### v3

@@ Abstract.
The main drawback of using generative AI models for advanced mathematics is that these models are not primarily logical reasoning engines. However, Large Language Models, and their refinements, can pick up on patterns in higher mathematics that are difficult for humans to see. By putting the design of generative AI models to their advantage, mathematicians may use them as powerful interactive assistants that can carry out laborious tasks, generate and debug code, check examples, formulate conjectures and more. We discuss how generative AI models can be used to advance mathematics research. We also discuss their integration with neuro-symbolic solvers, Computer Algebra Systems and formal proof assistants such as Lean.
@@ 1. Introduction
This limitation is inherent to the architecture of LLMs. They are fundamentally statistical, not logical, engines. As next-word predictors, they lack a built-in engine for formal symbolic deduction, meaning their emergent ‘reasoning’, called Natural Language Reasoning, is derived from statistical patterns in their training data.
@@
The AI landscape is rapidly evolving beyond standard LLMs, with Large Reasoning Models (LRMs) and Large Context Models (LCMs) emerging as the next generation of this technology. While built on the foundation of LLMs, these models have distinct architectures and methods of training. In particular, LRMs, such as Gemini, use hybrid neuro-symbolic systems which combine text generation with verification (see Subsection 3.8).
In Section 2, we describe the capabilities and limitations of LLMs and the nature of their training data. In Section 3, we explore the mechanics of generative AI, from the management of context windows to the geometric representation of mathematical language within high-dimensional embedding spaces. We also discuss the evolution of architecture from standard language models to Large Reasoning Models (LRMs) that utilize neuro-symbolic verification.
In Sections 4 through  6, we examine methods for influencing the outputs of these models. We discuss how they simulate computational environments and how prompt engineering can impose logical order on probabilistic outputs (Section 5).
In Section 7, we discuss how generative AI models are useful for questions in Combinatorial Group Theory. In Section 8 we discuss the integration of LLMs with Computer Algebra Systems (CAS) and formal proof assistants such as Lean.
Finally, in Section 9, we imagine a collaborative research partnership where a mathematician engages in an iterative dialogue with AI to discover new ideas and research directions.
2. Generative AI models
2.1. Mathematical training data for generative AI models
2.2. How much do generative AI models ‘know’?
In contrast, Symbolic Deduction is manipulation based on formal rules applied to abstract symbols, as in the operation of a Computer Algebra System. Current LLMs are built to imitate reasoning based on patterns, but are incapable of performing true symbolic deduction.
The significant advancement in reasoning capability has been achieved by neuro-symbolic systems that use an advanced form of Chain-of-Thought and Tree-of-Thought reasoning with a feedback loop for verification.
2.3. Drawbacks of using LLMs in mathematics
The primary drawback of using LLMs for mathematics is that they are probabilistic pattern-matchers, not logical reasoning engines. Their core function is to predict the next most likely word or symbol in a sequence, not to apply deterministic mathematical rules. Generative AI models are experts at reconstructing what a proof should ‘look like’. This limitation leads to several key problems:
3. Foundations of Generative AI
@@ 3.1. Tokens: the building blocks of language
3.2. The LLM as a probabilistic knowledge graph
Figure 1. A simplified view of an LLM as a ‘knowledge graph’.
3.3. Generative AI models and patterns in higher mathematics
A generative AI model can identify mathematical patterns that are not obvious to humans. It performs massive scale statistical analysis on the symbolic structure of mathematics itself. This allows it to simulate a mathematician’s intuition. It can find cross-disciplinary correlations and potential analogies that humans may not think of.
There are advantages of this capability. A generative AI model sees mathematics as a language with a complex grammar and vocabulary composed of symbols and diagrams. It learns the statistical relationships between the tokens in this language.
3.4. The embedding space for an LLM
Figure 2. A hypothetical embedding for $d=2$.
Generative AI models don’t ‘see’ symbols and equations. There is a map $\varphi:V\to\mathbb{R}^{d}$ (for $d\leq 13,000$ for large models) from the vocabulary $V$ of tokens to a high-dimensional vector space (the embedding space). The dimension $d$ is a model parameter. Relationships in the vocabulary are translated into geometric relationships in $\mathbb{R}^{d}$ [mikolov2013].
The model can detect geometric relationships in $\mathbb{R}^{d}$ that are difficult for humans to visualize for $d>>0$ and that are not obvious from the symbolic definitions.
3.5. The path of a word $w$ through an LLM
Let $\varphi(w)$ denote the initial embedding of the token word $w$ in $\mathbb{R}^{d}$. Its ‘meaning’ is its relationship to all other images of tokens in the space and their locations. Furthermore, in advanced contextual models, this meaning changes with context.
The model uses linear transformations $T_{q}$, $T_{k}$ and $T_{v}$ to project each token embedding $\varphi(w)\in\mathbb{R}^{d}$ into three specialized vectors which have different purposes: a query vector $q$ representing ‘what I am looking for’, a key vector $k$ encoding ‘what topic I represent’ and a value vector $v$ representing ‘what information I contain’. The maps
$T_{q},\quad T_{k},\quad T_{v}:\mathbb{R}^{d}\to\mathbb{R}^{d},$
project $\varphi(w)$ to three specialized vectors:
$q=T_{q}\varphi(w),\qquad k=T_{k}\varphi(w),\qquad v=T_{v}\varphi(w).$
The model determines relevancy scores by measuring the dot product between the query vector $q$ of a token and the key vectors $k_{i}$ of every token in the sequence. If $q$ and some $k_{i}$ point in similar directions, they obtain a high relevancy score.
It then uses these scores to create a new, context-rich vector as a weighted average of the value vectors. This new vector is then fed through a piecewise linear map to produce the output vector, $\zeta^{(\ell)}(w)$ for level $\ell$ of the model.
The process is repeated through many layers of the model; the output $\zeta^{(\ell)}(w)$ becomes the input for layer $\ell+1$. The final contextualized representation, $\zeta^{(N)}(w)$ is the output of the final layer $N$ and is different for every sentence it appears in.
The model generates an output sentence token-by-token in a loop. To predict the next token in the output, it uses the entire output sequence generated so far as its input. It then takes the final contextualized representation $\zeta^{(N)}(w_{s})$ of the last token $w_{s}$ from that sequence, passes it through one final linear map, and adds the most probable next token to the output sequence. This becomes the input for the next step.
Token Embedding $\varphi(w)\in\mathbb{R}^{d}$ Embedding space $\mathbb{R}^{d}$$q$ Query: ‘what I am looking for’ $k$ Key: ‘what topic I represent’ $v$ Value: ‘what information I contain’ $T_{q}$$T_{k}linear transformation$$T_{v}$linear transformationLayer $\ell$ output $\zeta^{(\ell)}(w)$ Stack of transformer blocks:
(attention + feed-forward
neural networks)
piecewise-linear maps,
repeated NN times
Final contextualized representation $\zeta^{(N)}(w)$
Figure 3. Projection of a token embedding $\varphi(w)$ in the embedding space $\mathbb{R}^{d}$ into specialized vectors ($q,k,v$) via linear transformations $T_{q},T_{k},T_{v}$, generation of the layer-$\ell$ output $\zeta^{(\ell)}(w)$, and repeated processing through a stack of transformer (neural network) blocks, including the feed-forward piecewise-linear map, producing the final contextualized representation $\zeta^{(N)}(w)$.
Token Embedding $\varphi(w)\in\mathbb{R}^{d}$ Embedding space $\mathbb{R}^{d}$$q$ Query: ‘what algorithm solves this?’ $k$ Key: ‘what topic I represent’ $v$ Value: ‘what information I contain’ $T_{q}$$T_{k}linear transformation$$T_{v}$linear transformationLayer $\ell$ output $\zeta^{(\ell)}(w)$ Stack of transformer blocks:
Chain-of-Thought generation
Python Script / PlanGenerateExternal Python interpreterExecute Deterministic logic
Computed resultFinal verified response
Neuro-symbolicfeedback loop
Figure 4. Unlike standard LLMs which predict tokens directly from the transformer stack, an LRM generates a computational plan (Python script), executes it in an external deterministic sandbox, then integrates the result back into the neural stream for the final response. Verification is built-in. The model no longer predicts mathematical answers. It predicts the algorithms needed to find the answers and it executes the algorithms.
3.6. Structure of the point cloud of token embeddings in an LLM
Let
$T^{(0)}:=\{\varphi(w)\mid w\ \text{a token}\}\subset\mathbb{R}^{d}$
denote the finite point cloud of embedded tokens. The Manifold Hypothesis of Machine Learning [TSL2000] asserts the existence of a smooth $n$-dimensional embedded submanifold
$M^{n}\subset\mathbb{R}^{d}$
such that the probability distribution supported on $T^{(0)}$ sits on a neighborhood of $M$.
In [MDC2025], the authors find that points in $T^{(0)}$ that are close in Euclidean distance exhibit ‘incompatible apparent dimensions’. That is, one point may have many nearby neighbors in $T^{(0)}$, while the other has far fewer.
This cannot occur if the point cloud is sampled from (or near) a single smooth embedded manifold $M^{n}\subset\mathbb{R}^{d}$. They conclude that there exists no low-dimensional smooth manifold $M$ in $\mathbb{R}^{d}$ whose neighborhoods model the local statistical structure of the token embedding point cloud. That is, the Manifold Hypothesis is violated.
Each transformer layer in the LLM implements a map
$f_{\ell}:\mathbb{R}^{d}\to\mathbb{R}^{d}$
built from linear projections (including $T_{q}$, $T_{k}$, $T_{v}$), attention mixing, and a piecewise-linear feed-forward network.
Define the point cloud at layer $\ell$ to be
$T^{(\ell)}:=f_{\ell}\circ f_{\ell-1}\circ\cdots\circ f_{1}(T^{(0)}).$
The observed behavior is compatible with the point clouds $T^{(\ell)}$ sampling neighborhoods of a stratified subset
$S=\bigsqcup_{i}S_{i}\subset\mathbb{R}^{d},$
where the strata $S_{i}$ have different dimensions and meet along singular loci. The singular points are those points where the space is not smooth or locally Euclidean. Near such points, no Euclidean ball intersects $S$ in a way resembling an open ball in $\mathbb{R}^{n}$ for any $n$.
The authors show that the maps $f_{\ell}$ preserve the types of singularities already present in $T^{(0)}$. The singularities correspond to polysemous words (words with multiple meanings). The different semantic regions of the embedded point cloud corresponding to polysemous words are ‘pinched’ together [JGZ2020]. Near the singularities, the LLM is unstable and ‘unusually sensitive’, leading to unexpected associations and incoherent reasoning. This may be one of several causes of model hallucinations.
[…]
Figure 5. On the left, (in blue) a smooth 1-dimensional manifold embedded in $\mathbb{R}^{2}$, where every point has a neighborhood homeomorphic to an open interval. On the right, (in red) a stratified subset of $\mathbb{R}^{2}$, hypothetically resembling an LLM token embedding space. It is a union of 1-dimensional strata crossing at a singular point and an isolated 0-dimensional stratum.
3.7. A single chat is a ‘context window’
@@
A generative AI model’s short-term memory is its context window, which consists of a single chat session. This window contains the entire conversation history, including uploaded documents and any corrections made by the user. Its capacity is measured in tokens: most standard models offer a 120K window which can hold a novel. Models like Gemini offer an unprecedented 1 million token input limit.
@@
3.8. Large Reasoning Models and Large Context Models
LLMs are predictive and fast. They ‘guess’ the end result but they may not do all the intermediate work.
Large Reasoning Models (LRMs), are a class of LLMs trained to carry out tasks that require multi-step deduction, critical thinking, and structured problem-solving. LRMs deliberate, explore various solution paths, evaluate intermediate steps and revise their reasoning.
LRM models integrate frameworks that support reasoning structures, such as ‘Chain-of-Thought’ and ‘Tree-of-Thought’ structures into their systems. We give some examples in Section 5.
@@ 3.8. Large Reasoning Models and Large Context Models
LRM models also have self-correcting capabilities. Their basis is a neuro-symbolic system which combines text generation with verification (see Figure 4). They have a feature known as ‘self-awareness’. LRMs are able to distinguish between requests that are language-based or computational. They make ‘decisions’ about delegating tasks to external tools such as a Python interpreter. They don’t try to predict mathematical answers, they predict the algorithms needed to find the answers, then they execute the algorithms and report back with the result.
@@
Google’s Gemini models demonstrate both LRM and LCM characteristics. These models also use internal ‘thinking processes’ for improved reasoning.
ChatGPT has a ‘reasoning mode’ in its most recent model. Reinforcement learning was used to teach the model to ‘think’ before generating answers, using what OpenAI refers to as ‘private chain-of-thought’ processes built into their system prompts. This allows the model to plan ahead and reason through tasks, performing a series of intermediate reasoning steps to assist in solving the problem.
Claude models use ‘extended thinking’ or the ability to ‘think out loud’.
These architectural advances are complemented by the practical integration of LLMs with symbolic computation tools.
4. Computational Capabilities
4.1. AI models simulate Computer Algebra Systems
@@
For example, the LLM may hallucinate functions that don’t exist in the Python libraries. In addition, the queries may give incorrect answers if the problems are inaccurately posed. The LLM could misunderstand the prompt, have bugs in its code or fail to handle boundary cases. Debugging AI-generated Python code may be needed.
@@
4.2. Comparison of models
For intermediate mathematical tasks, Gemini, ChatGPT, and Claude are all roughly in the same class with respect to performance. Though Gemini models performed slightly better in the standard benchmarks for competition-level mathematics, testing symbolic reasoning in algebra, calculus and number theory in 2024.
ChatGPT integrates features like WolframAlpha, which is a powerful Computer Algebra System, which can be enabled in the professional version. However, WolframAlpha is not designed for multi-step algorithms and it times-out on long tasks. Also, WolframAlpha is not optimized for integers. It makes errors with integers larger than $10^{15}$.
On the other hand, Python interpreters, available in Gemini and ChatGPT (see Subsection 4.1), have ‘arbitrary-precision’ for integers. They automatically adjust the memory used to store an integer as its value grows. We note that Python is not enabled in WolframAlpha.
Another ChatGPT feature called SciSpace has access to hundreds of millions of peer-reviewed journal papers.
5. Prompt engineering for mathematics
5.1. How the prompt shapes an LLM’s output
Structured prompts impose a logical framework on an LLM’s probabilistic generation process. The examples of prompts in the following subsections have been shown to be useful in guiding generative AI models on how to approach complex mathematical questions [kojima2022]. These methods are also used internally in advanced LRM models like Gemini and ChatGPT (in thinking mode).
5.2. $n$-shot prompts
An $n$-shot prompt provides the LLM with $n$ examples of the task that you would like it to perform.
Example: Zero-shot prompt
Determine the arithmetic progression corresponding to primes of the form $4n+1$ with last digit $1$, (such as $41,61,101,\dots$).
Example: One-shot prompt
Given that primes of the form $4n+1$ ending in $1$ (such as $41,61,101,\dots$) correspond to the arithmetic progression 20k+1{20k+1}, determine the arithmetic progression corresponding to primes of the form $4n+1$ with last digit $3$ (such as $13,53,73,113,\dots$).
Giving the LLM an example makes it more likely to give a correct answer.
5.3. Chain-of-Thought reasoning
This refers to the LLM thinking step-by-step, generating intermediate reasoning before arriving at the final answer [IBMToT]. If the LLM checks its work, this reduces errors.
Example prompt:
Your task is to generate a detailed description of the representation of the Weyl group of $\mathfrak{sl}_{3}(\mathbb{C})$ on the dual space $\mathfrak{h}^{*}$ of its Cartan subalgebra $\mathfrak{h}$.
Start by describing the Coxeter presentation of the group $W(A_{2})\cong S_{3}$ which has generators $s_{1},s_{2}$ and relations $s_{i}^{2}=e$ and $(s_{1}s_{2})^{3}=e$. Next, define the vector space $\mathfrak{h}^{*}$ and specify its basis of simple roots, $\alpha_{1},\alpha_{2}$.
Then define the representation $\rho$ of $W(A_{2})$ using the fact that the generators act as reflections on $\mathfrak{h}^{*}$.
Use the Weyl reflection formula to explicitly derive the matrices for $\rho(s_{1})$ and $\rho(s_{2})$ in the basis for $\mathfrak{h}^{*}$.
With this information, the next step is to verify that this is a valid representation. Please do this by performing the necessary matrix calculations to show that these matrices satisfy the defining relations of the group $W(A_{2})\cong S_{3}$.
An LRM may carry out internal reasoning as in Figure 6. A LLM can be explicitly prompted to use Chain-of-Thought reasoning.
[…]
Figure 6. Chain-of-thought diagram.
5.4. Tree-of-Thought reasoning
This refers to an AI model exploring multiple different deduction paths simultaneously, such that it evaluates its own progress at each step and pursues the most promising path [IBMToT].
Example prompt:
Let $p$ be a prime number, written $p=a_{n}a_{n-1}\dots a_{1}a_{0}$ in terms of its digits in base 10. Suppose that $n>1$, that is, suppose that $p$ has at least two digits. Your task is to determine the possible values of the last digit $a_{0}$.
An LRM may carry out internal reasoning as in Figure 7. A LLM can be explicitly prompted to use Tree-of-Thought reasoning.
Determine last digits of primes $p>10$ of the form $4n+1$ or $4n+3$ Primes $p>10$ cannot end in 0, 2, 4, 5, 6, 8. Candidate set for last digit is $\{1,3,7,9\}$ For $p=4n+1$ $10k+1\equiv 1\pmod{4}\implies k$ is even. Ex: 41 $10k+3\equiv 1\pmod{4}\implies k$ is odd. Ex: 13 $10k+7\equiv 1\pmod{4}\implies k$ is odd. Ex: 17 $10k+9\equiv 1\pmod{4}\implies k$ is even. Ex: 29 For $p=4n+3$ $10k+1\equiv 3\pmod{4}\implies k$ is odd. Ex: 11 $10k+3\equiv 3\pmod{4}\implies k$ is even. Ex: 23 $10k+7\equiv 3\pmod{4}\implies k$ is even. Ex: 47 $10k+9\equiv 3\pmod{4}\implies k$ is odd. Ex: 19 Conclusion: Both forms can end in any digit from the set $\{1,3,7,9\}$.
Figure 7. A Tree-of-Thought diagram for determining the possible last digits of primes $p>10$.
5.5. Using LLMs to generate Python code to run in SageMath